the sum of all five numbers is 30. the sum of the 5th number and 3rd number is 14. the 1st number is 1< twice the 2nd number. the 4th number 1> the 2nd. the 2nd and the 3rd number is 10.what are the five numbers??
Answer by checkley77(12844)

A+B+C+D+E=30
C+E=14 OR E=14-C
A=2B-1
D=B+1
B+C=10
2B-1B+10-B+1+14-C=30
3B-C=30-24
3B-C=6
B+C=10 ADD
-----------------
4B=16
B=16/4
B=4 ANS.
A=2*4-1
A=8-1
A=7 ANS.
D=4+1
D=5 ANS.
4+C=10
C=10-4
C=6 ANS.
6+E=14
E=14-6
C=8 ANS.
7+4+6+5+8=30
30=30